Consider the following digital circuit fi a b f2 с i B bc Do a fa b The following are equivalent expressions except (selec

iD bc c d- b- The following are comivalent expressions ex Select] in f2=2m(0,2,3,4,7) f2=TM(1,5,7) f2=bcl + bc + ab f2=b(

bc a b The following are equivalent expressions excep in every case): f1 [ Select ] f3=a(b+c)+b(a+c) f3=abc+b(a+c) f3=
Answer #1

Answers

f1=bc + ab +ab

f2=bcl + bc + alb

f3=abc + ab + bc

Explanation

f1

(a+c')b' = ab' + b'c'
a'b

f2

a'b
(b'+c)(b+c') = b'c' + bc

f3

a'b'c
(a+c')b = ab + bc'

please up vote